If Qd=20,000-3P Qs=15,000+2P. Calculate equilibrium price and quantity.

Calculate price elasticity of supply using point elasticity method. Also interpret the result.



1
Expert's answer
2021-12-05T18:58:21-0500

Solution:

At equilibrium: Qd = Qs

20,000 – 3P = 15,000 + 2P

20,000 – 15,000 = 2P + 3P

5,000 = 5P

P = 1,000

Equilibrium price = 1,000

Substitute in either the demand or supply function to derive equilibrium quantity:

Qd = 20,000 – 3P = 20,000 – 3(1,000) = 20,000 – 3,000 = 17,000

Equilibrium quantity = 17,000 units

 

Price elasticity of supply using point method:

PES = "\frac{\triangle Qs}{\triangle p}\times \frac{P}{Qs}"


"\frac{\triangle Qs}{\triangle p} = 2"


PES = "2\times\frac{1,000}{17,000} = 2 \times0.00588 = 0.01176"

Price elasticity of supply = 0.012


The price elasticity of supply is less than one, therefore, it is price inelastic.